# Life on a closed timelike curve

### v1

@@ V Conclusions
We stress that the present article does not argue in favor or against the existence of closed timelike curves. In fact, a proper discussion of the chronology protection conjecture [33, 34] in Gödel-type universes would require that we express the operator $J$ (defined in section II.2) in terms of quantum fields, and that we study the backreaction of the vacuum fluctuations on the spacetime metric [35], which is beyond the scope of this article [^2] Indeed, CTCs are known to generate pathologies in interacting quantum field theories (see [36] for a review). The reader can see [37, 38, 39] for derivations of the chronology protection conjecture within string theory and holography.. Rather, the take-home message of the present article is that, in a hypothetical universe where closed timelike curves were possible, time travel would not take place in the form that is usually depicted in science fiction literature. As it often happens, Nature is more creative than us.
Finally, we remark that our main results remain valid in an arbitrary background spacetime, provided that the CTC of interest is the orbit of a periodic one-parameter family of symmetries of the metric. Here, we chose Gödel’s metric because it describes a particularly simple universe, and it played a significant role in the historical development of the discussion around CTCs. Gödel’s metric is also a convenient spacetime in that it does not possess singularities, which may otherwise cause trouble when one tries to formalize the quantum theory.
@@ Acknowledgements
This work was supported by a Vanderbilt’s Seeding Success Grant.
Appendix A The Hamiltonian of the spaceship
Fix an observable scalar field of interest, $F(x^{\mu})$, and evaluate it along the CTC (3). The result is a one-parameter family of Hermitian operators $F(\phi)$ that an observer can measure during their journey. Working in the Heisenberg picture (which is the fundamental picture in relativity, where no preferred time is given a priori), the action of the operator $J$ on $F$ is determined by the following identity (see [15, §2.4], [40, §3.2], and [41, §2]):
@@ Appendix A The Hamiltonian of the spaceship
This is the defining property of the unitary group (5), since it makes manifest the link between the operators $U(b)$ (left-hand side) and the spacetime symmetries $\mathcal{G}_{b}$ (right-hand side). Averaging the above equation over a generic quantum state |$\ket{\Psi}$, and setting $\phi=0$, we obtain
@@
Appendix B Time gradients inside the spaceship
@@
The following argument is a straightforward adaptation of the derivation of the gravitational redshift formula in stationary spacetimes, see e.g. [42, §5.3].
@@
We assume that the support of the spaceship is contained within the spacetime region where $\partial_{\phi}$ is timelike. Suppose that there are two passengers, Alice and Bob, who take two different seats in the spaceship. Alice sits on the worldline $\gamma_{A}(\phi)=(t_{A},\phi,r_{A},z_{A})$, and Bob sits on the worldline $\gamma_{B}(\phi)=(t_{B},\phi,r_{B},z_{B})$. Alice sends two light pulses to Bob, the first time when she is on the event $(t_{A},\phi_{A1},r_{A},z_{A})$, and the second time when she is on the event $(t_{A},\phi_{A2},r_{A},z_{A})$. The first lightpulse travels along a null geodesic $\gamma_{1}$ that begins at $(t_{A},\phi_{A1},r_{A},z_{A})$ and terminates on Bob’s worldline, at some event $(t_{B},\phi_{B1},r_{B},t_{B})$. Now, we note that $\gamma_{1}$ is a null geodesic if and only if $\mathcal{G}_{b}\gamma_{1}$ is also a null geodesic, for any $b$ [43, §25.2]. Therefore, since the second light pulse is generated in the event $(t_{A},\phi_{A2},r_{A},z_{A})$, and points in the same direction as the first, it will follow the null geodesic $\gamma_{2}=\mathcal{G}_{\phi_{A2}-\phi_{A1}}\gamma_{1}$, thereby reaching Bob in the event

### v2

@@ V Conclusions
We stress that the present article does not argue in favor or against the existence of closed timelike curves. In fact, a proper discussion of the chronology protection conjecture [41, 42] in Gödel-type universes would require that we express the operator $J$ (defined in section II.2) in terms of quantum fields, and that we study the backreaction of the vacuum fluctuations on the spacetime metric [43, 44, 45], which is beyond the scope of this article [^5] Indeed, CTCs are known to generate pathologies in interacting quantum field theories (see [46] for a review). The reader can see [47, 48, 49] for derivations of the chronology protection conjecture within string theory and holography.. Rather, the take-home message is that, in a hypothetical universe with CTCs, time travel would not take place in the form that is usually depicted in science fiction. In fact, on CTCs, thermal fluctuations destroy macroscopic causation, and erase all memories (as conjectured in [50]). As it often happens, Nature is more creative than us.
Finally, we stress again that our main results are valid in an arbitrary background spacetime (including charged Kerr black holes [51, §12.3]), provided that the CTC of interest is the orbit of a periodic one-parameter family of symmetries of the metric. This happens in all axisymmetric models whose rotation Killing field $\partial_{\phi}$ becomes timelike somewhere.
@@ Acknowledgements
This work was supported by a Vanderbilt’s Seeding Success Grant. We thank all referees for the useful input.
Appendix A A more “in depth” proof of (6)
Let us first recall what a symmetry is, in the words of Steven Weinberg [22, §2.2]:
A symmetry transformation is a change in our point of view that does not change the result of possible experiments. If an observer $O$ sees a system in a state represented by a ray $\mathscr{R}$ or $\mathscr{R}_{1}$ or $\mathscr{R}_{2}$…, then an equivalent observer $O^{\prime}$ who looks at the same system will observe it in a different state, represented by a ray $\mathscr{R}^{\prime}$ or $\mathscr{R}^{\prime}_{1}$ or $\mathscr{R}^{\prime}_{2}$…, respectively, but the two observers must find the same probabilities
$P(\mathscr{R}\rightarrow\mathscr{R}_{n})=P(\mathscr{R}^{\prime}\rightarrow\mathscr{R}^{\prime}_{n})\,.$ (24)
In our case, we are considering the spacetime symmetry $\phi^{\prime}=\phi-b$. This corresponds to a change of perspective from an observer $O$ located at $\phi{=}0$ to an equivalent observer $O^{\prime}$ located at $\phi{=}b$. In the case where $b{=}2\pi$, the second observer is located at $\phi{=}2\pi$, which is the same as $\phi{=}0$. Hence, there is no change of perspective in this case, because the observers $O$ and $O^{\prime}$ coincide. Thus, we have that $\mathscr{R}=\mathscr{R}^{\prime}$ and $\mathscr{R}_{n}=\mathscr{R}^{\prime}_{n}$, and the identity $e^{i2\pi J}=\text{const}=e^{i\alpha}$ follows immediately from the definition of ray and the ability to make superpositions.
A.1 Parallel realities are incompatible with relativity…
Looking at the above proof, the careful reader may spot the following potential loophole: What if we treat the observers $O$ and $O^{\prime}$ as genuinely different, by assigning to them a “winding number” of the phase $\phi$? In this case, there might be two “parallel” quantum evolutions (and two parallel realities) related to the fact that $O$ and $O^{\prime}$ are formally occupying the same location $\phi{=}0$ in spacetime, but carry a different winding number, and thus observe the system to be in a different quantum state. Wouldn’t this invalidate the proof?
Such a scenario does not fit within standard relativity, as it would imply that two observers at the same spacetime point can disagree on the value of a Lorentz scalar at that point. It also invalidates the way we conceptualize the spacetime manifold, since according to Einstein himself a spacetime point (or event) is uniquely identified by the list of all “facts” that happen in it [52], and two observes cannot disagree on such facts. About this, Misner, Thorne & Wheeler [53, §1.2] wrote:
Happily, nature provides its own way to localize a point in spacetime, as Einstein was the first to emphasize. Characterize the point by what happens there! Give a point in spacetime the name “event”. Where the event lies is defined as clearly and sharply as where two straws cross each other in a barn full of hay.
A.2 …or not?
While the above argument rules out parallel realities at the same event within standard physics, it also allows us to define a more conservative notion of “parallel realities” that fits perfectly well within canonical relativity.
Suppose that $O$ is located at $\phi=0$ and $O^{\prime}$ is located at $\phi=2\pi$. Assume that $O$ measures, say, the Faraday scalar $F^{\mu\nu}F_{\mu\nu}$, and gets $5$, while $O^{\prime}$ measures the same scalar $F^{\mu^{\prime}\nu^{\prime}}F_{\mu^{\prime}\nu^{\prime}}$, and gets $-3$. Then, $O$ and $O^{\prime}$ are not occupying the same event in spacetime. Instead, we should interpret the disagreement between $O$ and $O^{\prime}$ as a signal that $\{\phi=0\}\not\equiv\{\phi=2\pi\}$, and the spacetime manifold is actually a multi-sheeted Riemann-type surface, where the winding number $\left\lfloor\phi/(2\pi)\right\rfloor$ determines which sheet an observer is on. As a result, the geometry of the system is now very different from the one considered in section II.1, in that:
  • (a)
    The spacetime is no longer a Gödel-type universe, since it possesses multiple $\phi-$sheets, and consequently a conical singularity at $r=0$.
  • (b)
    The curve (3) is no longer a CTC, since it never goes back to where it started. Instead, it just “winds” around the Riemann-type surface like a helix, with ever-increasing $\phi$.
In conclusion, the parallel-reality scenario where $O{\not\equiv}O^{\prime}$ is formally allowed in relativity, but it corresponds to a geometry without CTCs. Clearly, this is not the type of spacetime considered in the present article.
Appendix B The Hamiltonian of the spaceship
Fix an observable scalar field of interest, $F(x^{\mu})$, and evaluate it along the CTC (3). The result is a one-parameter family of Hermitian operators $F(\phi)$ that an observer can measure during their journey. Working in the Heisenberg picture (which is the fundamental picture in relativity, where no preferred time is given a priori), the action of the operator $J$ on $F$ is determined by the following identity (see [22, §2.4], [54, §3.2], and [55, §2]):
@@ Appendix B The Hamiltonian of the spaceship
This is the defining property of the unitary group (5), since it makes manifest the link between the operators $U(b)$ (left-hand side) and the spacetime symmetries $\mathcal{G}_{b}$ (right-hand side) [^6] Note that (25) can directly be used to derive (6). In fact, if we set $b=2\pi$ in (25), and use the periodicity of $\phi$ together with the single-valuedness of fields, we obtain $e^{-i2\pi J}F(\phi)e^{i2\pi J}=F(\phi)$. Thus, $e^{i2\pi J}$ must commute with all local observables. If we assume that all physical observables are built out of local ones (which is the case for quantum field theory), we conclude that $e^{i2\pi J}$ must be proportional to the identity, giving again (6).. Averaging the above equation over a generic quantum state |$\ket{\Psi}$, and setting $\phi=0$, we obtain
@@
Appendix C Finite-size corrections
In this Appendix, we expand the discussion of II.5.
C.1 Scale ambiguity of Hamiltonians in General Relativity
Consider a generic spacetime manifold, with a Killing vector field $\mathcal{K}$ that is locally timelike. Then, there exists a local coordinate system $\{x^{0},x^{1},x^{2},x^{3}\}$ such that $\mathcal{K}=\partial_{0}$, where $x^{0}$ is a local time coordinate. Correspondingly, we can define a Hamiltonian $H$ such that $e^{-iH\theta}$ is the unitary operator representing the translation $x^{0}\rightarrow x^{0}+\theta$. On the other hand, if $\mathcal{K}$ is a Killing vector, then also $\tilde{\mathcal{K}}=a\mathcal{K}$ is a Killing vector, for any spacetime-independent constant $a{>}0$. Therefore, we can define a new coordinate system $\{\tilde{x}^{0},\tilde{x}^{1},\tilde{x}^{2},\tilde{x}^{3}\}{=}\{x^{0}/a,x^{1},x^{2},x^{3}\}$ such that $\partial_{\tilde{0}}=a\partial_{0}=a\mathcal{K}=\tilde{\mathcal{K}}$. The associated Hamiltonian operator $\tilde{H}$ will then generate the transformations $e^{-i\tilde{H}\tilde{\theta}}$, which represent the translations $\tilde{x}^{0}\rightarrow\tilde{x}^{0}+\tilde{\theta}$. But since $\tilde{x}^{0}=x^{0}/a$, the unitary operator $e^{-i\tilde{H}\tilde{\theta}}$ must coincide with the unitary operator $e^{-iHa\tilde{\theta}}$ (for all $\tilde{\theta}$), so that $\tilde{H}=aH$. We conclude that, in general relativity, there is an intrinsic ambiguity in how to define the Hamiltonian operator, related to our ability to rescale the time coordinate [^7] In the non-relativistic limit, this ambiguity becomes the freedom of adding an arbitrary constant to the gravitational potential. Here is the proof. Write $H=m{+}H_{\text{nr}}$ ($m$ is the rest mass, and $H_{\text{nr}}$ is the non-relativistic Hamiltonian) and $a=1{+}\epsilon$. Then, $\tilde{H}=m+H_{\text{nr}}+\epsilon m+\epsilon H_{\text{nr}}$. In the non-relativistic limit, $\epsilon$ and $H_{\text{nr}}$ tend to zero, and the term $\epsilon H_{\text{nr}}$ may be neglected, being quadratic in infinitesimals. Therefore, we can write $\tilde{H}\approx m+\tilde{H}_{\text{nr}}$, with $\tilde{H}_{\text{nr}}=H_{\text{nr}}+m\epsilon$, where the latter term can be reabsorbed in the definition of the gravitational potential.. This ambiguity is expressed by the following group of redefinitions:
$\begin{cases}\tilde{\mathcal{K}}=a\mathcal{K}\,,\\ \tilde{H}=aH\,,\\ \tilde{x}^{0}=x^{0}/a\,.\end{cases}$ (28)
So, how do we choose the “most natural” value of $a$? We would like our time coordinate to coincide with the proper time of a stationary observer. Equivalently, we would like to choose the scale $a$ such that $g(a\mathcal{K},a\mathcal{K}){=}{-}1$. However, in general, it may be impossible to enforce this identity simultaneously at all locations, since
$\begin{split}\partial_{\alpha}[g(a\mathcal{K},a\mathcal{K})]={}&a^{2}\nabla_{\alpha}(\mathcal{K}^{\mu}\mathcal{K}_{\mu})\\ ={}&2a^{2}\mathcal{K}^{\mu}\nabla_{\alpha}\mathcal{K}_{\mu}\\ ={}&-2a^{2}\mathcal{K}^{\mu}\nabla_{\mu}\mathcal{K}_{\alpha}\,,\\ \end{split}$ (29)
which does not vanish if the stationary observer is accelerating. Therefore, we need to choose a preferred stationary worldline $(x^{1}_{p},x^{2}_{p},x^{3}_{p})$, where $g(a\mathcal{K},a\mathcal{K})$ equals $-1$, and this uniquely specifies the value of $a$. When we make this choice, the time coordinate will agree with the proper time measured by the observer traveling on $(x^{1}_{p},x^{2}_{p},x^{3}_{p})$, but it may not be the proper time of the neighboring observers. Similarly, the associated Hamiltonian will quantify the energy measured by detectors located at $(x^{1}_{p},x^{2}_{p},x^{3}_{p})$, while in other locations it will be a redshifted (or blueshifted) energy.
In section II.3, we employed the ambiguity (28) to adjust our definition of $H$ in (8). In particular, we have chosen $a$ such that $g(a\partial_{\phi},a\partial_{\phi})=-1$ on the CTC of interest. In an extended object, one may set $g(a\partial_{\phi},a\partial_{\phi})=-1$ in the object’s “center” (in the proximity of a central clock), and there will be some redshift (or blueshift) near the boundaries. However, the spectrum of $H$ will still have the form $E_{n}=\text{const}\times n$, since $a$ is just a constant factor.
C.2 Gradients in the perceived time: A thought experiment
@@
The following argument is a straightforward adaptation of the derivation of the gravitational redshift formula in stationary spacetimes, see e.g. [56, §5.3].
@@
We assume that the support of the spaceship is contained within the spacetime region where $\partial_{\phi}$ is timelike. Suppose that there are two passengers, Alice and Bob, who take two different seats in the spaceship. Alice sits on the worldline $\gamma_{A}(\phi)=(t_{A},\phi,r_{A},z_{A})$, and Bob sits on the worldline $\gamma_{B}(\phi)=(t_{B},\phi,r_{B},z_{B})$. Alice sends two light pulses to Bob, the first time when she is on the event $(t_{A},\phi_{A1},r_{A},z_{A})$, and the second time when she is on the event $(t_{A},\phi_{A2},r_{A},z_{A})$. The first lightpulse travels along a null geodesic $\gamma_{1}$ that begins at $(t_{A},\phi_{A1},r_{A},z_{A})$ and terminates on Bob’s worldline, at some event $(t_{B},\phi_{B1},r_{B},t_{B})$. Now, we note that $\gamma_{1}$ is a null geodesic if and only if $\mathcal{G}_{b}\gamma_{1}$ is also a null geodesic, for any $b$ [53, §25.2]. Therefore, since the second light pulse is generated in the event $(t_{A},\phi_{A2},r_{A},z_{A})$, and points in the same direction as the first, it will follow the null geodesic $\gamma_{2}=\mathcal{G}_{\phi_{A2}-\phi_{A1}}\gamma_{1}$, thereby reaching Bob in the event